OpenStudy (anonymous):

What is the slope of the line that passes through (3, –7) and (–1, 1)? undefined 0 –2 2

OpenStudy (hba):

m=y2-y1/x2-x1

OpenStudy (anonymous):

\[\frac{ y-difference }{ x-difference }\]

OpenStudy (hba):

Use the formula

OpenStudy (hba):

Show me the working

OpenStudy (hba):

And I will tell u wheter ur right or wrong

hero (hero):

Hint: use y = mx + b for each point to get: -7 = 3m + b 1 = -m + b Isolate b in each equation to get: b = -3m - 7 b = 1 + m Set b = b to get: -3m - 7 = 1 + m Add 3m to both sides, then subtract 1 from both sides to to get: 4m = -8 Divide both sides by 4 to get m = -8/4 m = -2 Since m represents the slope of a line, the slope, m, is -2

OpenStudy (anonymous):

its m=-2 got it thanks

OpenStudy (hba):

You are welcome